OpenStudy (anonymous):

after 6 years, an investment is worth $8479. if the money was invested at 9% interest compounded semiannually what was the initial investment

OpenStudy (anonymous):

P is the principal (the initial amount you borrow or deposit) r is the annual rate of interest (percentage) n is the number of years the amount is deposited or borrowed for. A is the amount of money accumulated after n years, including interest. When the interest is compounded once a year: A = P(1 + r)^n 8479 = P (1+0.09/2)^12 8479 = P (1.69588) P = 4999.76